OpenStudy (anonymous):

What is the hypotenuse-leg theorem? A. If the hypotenuse of a right triangle is congruent to the corresponding part of another right triangle, then the triangles are congruent. B. If the hypotenuse and one leg of a right triangle are congruent to the corresponding parts of another right triangle, then the triangles are congruent. C. If the hypotenuse and one leg of a right triangle are similar to the corresponding parts of another right triangle, then the triangles are congruent.

OpenStudy (anonymous):

if the hypotenuse and leg of a right triangle are congruent to the hypotenuse and leg of another right triangle, then the triangles are congruent

OpenStudy (anonymous):

HI.. \[\LARGE \color{red}{\mathbb{WELCOME}}~~\color{orange}{\mathbb{TO}}~~\color{teal}{\mathbb{OPENSTUDY}}~\color{green}{!!}~~~(:\]